# Beyond Benchmark: LLMs Evaluation with an Anthropomorphic and Value-oriented Roadmap

### v1

@@ 1 Introduction
[…]
leaf/.style= my-box, minimum height=1.5em, fill=macaron-leaf, text=black, align=center, font=, inner xsep=2pt, inner ysep=4pt, line width=0.8pt, afterthought=, ,
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=center, text centered, minimum width=5em, edge+=darkgray,line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, , where level=0fill=none,draw=none,font=,rotate=90,anchor=north, where level=1fill=macaron-level1,font=, where level=2fill=macaron-level2, where level=3fill=macaron-level3, where level=4fill=macaron-level4, where level=5fill=macaron-level5, where level=1text width=16em,font=, where level=2text width=16em,font=, where level=3text width=19em,font=, where level=4text width=22em,font=, where level=5text width=18em,font=, [ [ 1. Introduction ] [ 2. LLM Evaluation Harness
/ Engineering
[ 2.1 Prominent LLM
Evaluation Harness, leaf [ Task Support, leaf ] [ Diversity of Objectives, leaf ] [ Usability, leaf ] [ Community Activity, leaf ] ] [ 2.2 Practical Guide of Modular
LLM Evaluation System, leaf [ Benchmark or Dataset Hub, leaf ] [ Model Hub, leaf ] [ Prompting Module, leaf ] [ Metrics Module, leaf ] [ Tasks Module, leaf ] [ Leaderboards and Arena Module, leaf ] [ Analysis Module, leaf ] ] ] [ 3. Anthropomorphic Evaluation
Taxonomy: IQ PQ and EQ
[ 3.1 Intelligence Quotient (IQ)
General Intelligence Evaluation, leaf ] [ 3.2 Professional Quotient (PQ)
Professional Expertise Evaluation, leaf [ Healthcare Domain, leaf ] [ Financial Domain, leaf ] [ Legal Domain, leaf ] [ Telecommunications Domain, leaf ] [ Coding Domain, leaf ] [ Software Domain, leaf ] [ Science Domain, leaf ] ] [ 3.3 Emotional Quotient (EQ)
Alignment Ability Evaluation, leaf ] ] [ 4. Value-Oriented Evaluation
(VQ - Value Quotient)
[ 4.1 Economic Value, leaf [ Cost-Benefit Ratio, leaf ] [ Return on Investment, leaf ] [ Productivity Improvement, leaf ] [ Market Acceptance, leaf ] ] [ 4.2 Social Value, leaf [ User Satisfaction, leaf ] [ Knowledge Dissemination Efficiency, leaf ] [ Public Service Improvement, leaf ] [ Education Quality Improvement, leaf ] ] [ 4.3 Ethical Value, leaf [ Fairness, leaf ] [ Transparency, leaf ] [ Privacy Protection, leaf ] [ Bias Detection, leaf ] ] [ 4.4 Environmental Value, leaf [ Energy Efficiency, leaf ] [ Carbon Footprint, leaf ] [ Sustainability, leaf ] ] ] [ 5. LLM System or
Application Evaluation
[ 5.1 RAG, leaf ] [ 5.2 Agent, leaf ] [ 5.3 Chatbot, leaf ] ] [ 6. Challenges & Future Perspectives [ 6.1 Enhanced Statistical Analysis, leaf ] [ 6.2 Composite Evaluation Systems, leaf ] [ 6.3 Interpretability and Explainability, leaf ] [ 6.4 User-Centric Benchmark, leaf ] [ 6.5 Human in the Loop Evaluation, leaf ] [ 6.6 Analytical Failure Exploration, leaf ] [ 6.7 Dynamic and Agentic Evaluation, leaf ] [ 6.8 Reproducibility Reliability Robustness, leaf ] ] [ 7. Conclusions ] ]
@@ 2.2.4 Metrics Module
Lexical and Morphological Metrics: These metrics focus on token- or character-level correspondence and morphological variation. Traditional n-gram overlap measures such as BLEU [39] and ROUGE-N/L [41] quantify the proportion of exact contiguous matches between hypothesis and reference, while edit-distance scores like Translation Error Rate (TER) gauge the minimal sequence of insertions, deletions, and substitutions required to transform one string into another. Complementing these, word-order distances—including RIBES [42] and Kendall’s $\tau$ —penalize token reordering, providing insight into the impact of syntactic shifts on surface similarity. For tasks sensitive to finer-grained discrepancies, character error rate (CER) and word error rate (WER) compute the frequencies of low-level insertion, deletion, and substitution errors, as commonly used in ASR and OCR evaluation. Subword-overlap metrics such as chrF [43] and BPE-F1 further refine this analysis by measuring similarity over character n-grams or byte-pair encoded segments, thus capturing partial matches that evade pure token-level statistics.
@@
Syntactic Metrics: it assesses the preservation of grammatical structure and targeted syntactic phenomena. Constituency and dependency parse-tree matching metrics—PARSEVAL [44] precision, recall, and F1 for bracket structures, alongside Unlabeled and Labeled Attachment Scores (UAS/LAS) [45] for dependency relations—offer a principled basis for comparing predicted and gold parses. To probe a model’s command of specific constructions, targeted syntactic evaluation frameworks such as Targeted Syntactic Evaluation (TSE) [46] deploy minimal-pair sentences to test capabilities like subject–verb agreement, while syntactic tree-edit distance measures the minimal sequence of tree operations to align two parse trees, yielding a granular account of structural divergence.
@@
Semantic Metrics: At semantic level, evaluation emphasizes meaning preservation, inference, and fidelity. Embedding-based approaches—BERTScore [47], MoverScore [48], BLEURT [49], COMET [50], and similar methods—leverage contextualized vectors extracted from pre-trained models to compute cosine similarities or Earth Mover’s distances, thus capturing nuanced semantic alignment between hypothesis and reference. Entailment-driven metrics such as the Document-Aware Entailment model (DAE) [51] treats the generation task as a natural language inference problem, classifying whether outputs are entailed by, neutral to, or contradictory with source texts. Question-answering frameworks, including QuestEval [52] and QAFactEval [53], automatically generate questions from the source or candidate summary and compare model-predicted answers to gold responses, thereby quantifying semantic fidelity via answer accuracy. Specialized LLM-based judges—either prompt-based few-shot evaluators or fine-tuned discrimination models—have emerged as learnable arbiters of output quality, scoring on dimensions such as factuality, coherence, and naturalness [54].

### v2

@@ 1 Introduction
1. Introduction 2. LLM Evaluation Harness / Engineering 2.1 Prominent LLM Evaluation Harness Task Support Diversity of Objectives Usability Community Activity 2.2 Practical Guide of Modular LLM Evaluation System Benchmark or Dataset Hub Model Hub Prompting Module Metrics Module Tasks Module Leaderboards and Arena Module Analysis Module 3. Anthropomorphic Evaluation Taxonomy: IQ PQ and EQ 3.1 Intelligence Quotient (IQ) General Intelligence Evaluation 3.2 Professional Quotient (PQ) Professional Expertise Evaluation Healthcare Domain Financial Domain Legal Domain Telecommunications Domain Coding Domain Software Domain Science Domain 3.3 Emotional Quotient (EQ) Alignment Ability Evaluation 4. Value-Oriented Evaluation (VQ - Value Quotient) 4.1 Economic Value Cost-Benefit Ratio Return on Investment Productivity Improvement Market Acceptance 4.2 Social Value User Satisfaction Knowledge Dissemination Efficiency Public Service Improvement Education Quality Improvement 4.3 Ethical Value Fairness Transparency Privacy Protection Bias Detection 4.4 Environmental Value Energy Efficiency Carbon Footprint Sustainability 5. LLM System or Application Evaluation 5.1 RAG 5.2 Agent 5.3 Chatbot 6. Challenges & Future Perspectives 6.1 Enhanced Statistical Analysis 6.2 Composite Evaluation Systems 6.3 Interpretability and Explainability 6.4 User-Centric Benchmark 6.5 Human in the Loop Evaluation 6.6 Analytical Failure Exploration 6.7 Dynamic and Agentic Evaluation 6.8 Reproducibility Reliability Robustness 7. Conclusions
@@ 2.2.4 Metrics Module
Lexical and Morphological Metrics: These metrics focus on token‐ or character‐level correspondence and morphological variation. Traditional n-gram overlap measures such as BLEU [39] and ROUGE-N/L [41] quantify the proportion of exact contiguous matches between hypothesis and reference, while edit-distance scores like Translation Error Rate (TER) gauge the minimal sequence of insertions, deletions, and substitutions required to transform one string into another. Complementing these, word-order distances—including RIBES [42] and Kendall’s $\tau$ —penalize token reordering, providing insight into the impact of syntactic shifts on surface similarity. For tasks sensitive to finer‐grained discrepancies, character error rate (CER) and word error rate (WER) compute the frequencies of low‐level insertion, deletion, and substitution errors, as commonly used in ASR and OCR evaluation. Subword‐overlap metrics such as chrF [43] and BPE-F1 further refine this analysis by measuring similarity over character n-grams or byte-pair encoded segments, thus capturing partial matches that evade pure token-level statistics.
@@
Syntactic Metrics: it assesses the preservation of grammatical structure and targeted syntactic phenomena. Constituency and dependency parse‐tree matching metrics—PARSEVAL [44] precision, recall, and F1 for bracket structures, alongside Unlabeled and Labeled Attachment Scores (UAS/LAS) [45] for dependency relations—offer a principled basis for comparing predicted and gold parses. To probe a model’s command of specific constructions, targeted syntactic evaluation frameworks such as Targeted Syntactic Evaluation (TSE) [46] deploy minimal‐pair sentences to test capabilities like subject–verb agreement, while syntactic tree-edit distance measures the minimal sequence of tree operations to align two parse trees, yielding a granular account of structural divergence.
@@
Semantic Metrics: At semantic level, evaluation emphasizes meaning preservation, inference, and fidelity. Embedding-based approaches—BERTScore [47], MoverScore [48], BLEURT [49], COMET [50], and similar methods—leverage contextualized vectors extracted from pre-trained models to compute cosine similarities or Earth Mover’s distances, thus capturing nuanced semantic alignment between hypothesis and reference. Entailment-driven metrics such as the Document‐Aware Entailment model (DAE) [51] treats the generation task as a natural language inference problem, classifying whether outputs are entailed by, neutral to, or contradictory with source texts. Question‐answering frameworks, including QuestEval [52] and QAFactEval [53], automatically generate questions from the source or candidate summary and compare model‐predicted answers to gold responses, thereby quantifying semantic fidelity via answer accuracy. Specialized LLM‐based judges—either prompt-based few-shot evaluators or fine-tuned discrimination models—have emerged as learnable arbiters of output quality, scoring on dimensions such as factuality, coherence, and naturalness [54].
@@ 3.2 Professional Expertise Evaluation (PQ)
Owl-Bench [157] Beihang University QA pairs, multiple-choice questions
9 distinct subdomains including information security
link
Software SWE-bench [158] Princeton NLP Real-world software problems from GitHub
Assesses ability to generate patches for described issues
link
OpsEval [140] Tsinghua University Wired network ops, 5G, database ops Evaluates proficiency in practical applications link